Grade 11Differential Calculus

find dy/dx when
y= e^(3x) sin 4x . 2^x

Sneha Dusaria
12 Years agoGrade 11

1 Answer

Arun Kumar
12 Years ago
Hello Student,
\\ \\\frac{d}{dx}\left(e^{\left(3x\right)}\sin \left(4x0.2^x\right)\right)\: \\=\:e^{3x}\left(3\sin \left(4\cdot \:5^{-x}x\right)-4\cdot \:5^{-x}\left(x\ln \left(5\right)-1\right)\cos \left(4\cdot \:5^{-x}x\right)\right)$ \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty